OpenStudy (anonymous):

Which is the slope and y-intercept of the following equation? y = 8x + 1

OpenStudy (anonymous):

Do you know slope y -intercept formula?

OpenStudy (anonymous):

no, unfortuanately.

OpenStudy (anonymous):

Please look it up, so you can learn!

OpenStudy (anonymous):

Paul is God. He got me through calculus, so I'm sure he can get you through algebra. Read through this whole section, you'll learn a lot. http://tutorial.math.lamar.edu/Classes/Alg/Lines.aspx

OpenStudy (anonymous):

Brinet comes to rescue you :)

OpenStudy (anonymous):

Chlor, you can still answer, I just wanted to pass on these notes. They are awesome.

OpenStudy (anonymous):

thanks ill read these and see if they help

OpenStudy (anonymous):

Just give me any idea about slope!

OpenStudy (anonymous):

@ewokkman1, your y-intercept will be when x=0. Look at any graph and you can understand this visually. So plug x=0 into your equation and see what you get.

OpenStudy (anonymous):

so the y-intercept will =9?

OpenStudy (anonymous):

What is 8*0?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

y = 8*0 +1

OpenStudy (anonymous):

oh...... i think i understand it. So my y-intercept would would equal 1

OpenStudy (anonymous):

Yup! :-)

OpenStudy (anonymous):

And if you ever have it where it's not in y=mx+b form, first step is to always always always get it in that form. No coefficient next to the y. Make sure y is by itself. You'll understand more when you read through those notes, they explain it really well.

OpenStudy (anonymous):

So what is "m" in your case?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Yep. And if you want to graph the line by starting at a given point, you want to move up 8 in the positive direction, and over 1 in the positive direction. They don't really specify this, but it's implied. m= 8/1 = 8. That helps if you don't feel like plugging in another point but just graphing the function. If it's -8, then that mean move down 8 in the negative direction and over 1 in the positive direction.